OpenStudy (anonymous):

help please Write the equation of the line that is perpendicular to the line y = three halvesx + 4 and passes through the point (-6, 3). y = three halvesx - 1 y = three halvesx + 3 y = -two thirdsx - 1 y = -two thirdsx + 3 Write the equation of the line that is parallel to the line y = -3x + 12 and passes through the point (-1, 6). y = one thirdx + 7 y = -3x + 3 y = one thirdx + 3 y = -3x + 7

OpenStudy (anonymous):

help please

OpenStudy (anonymous):

the perpendicular line has the opposite slope of the first one. The slope is the number in front of x: 3/2. Switch it around and you have the start of the perpendicular slope: 2/3. (So the first two answers are out already). Now also alter the sign: -2/3. Now you know the equation must look like y=-2/3 x + some number All you need to do now is check which of the last two equations includes the point (-6,3). Plug in -6 for x and 3 for y of the equation and let's call 'some number' b . \[3= - \frac{ 2 }{ 3 } . (-6) + b\] Now you calculate the value of -(2/3) . 6 and place it at the other side of the = sign (left of it), and you'll get a sum. When you calculate the sum you know what b must be. \[3=4 + b\] \[3-4=b\] -1=b So the equation should be y=-2/3 x -1 Now try to apply the same principal for the second question

OpenStudy (anonymous):

Only this time you must use the parallel equation, which means "same slope". So an equation with the same number in front of x as the given equation.

OpenStudy (anonymous):

so for the same slope you have to add both sides and find B ?

OpenStudy (anonymous):

for the same slope you just look at the original equation. What number is in front of x in y=-3x+12?

OpenStudy (anonymous):

so u do +3 +3 right

OpenStudy (anonymous):

i got D for it

OpenStudy (anonymous):

Look closer, what sign is in front of the 3 in the original equation?

OpenStudy (anonymous):

- sign

OpenStudy (anonymous):

yup, so the slope is -3

OpenStudy (anonymous):

so that would make it B

OpenStudy (anonymous):

which two equations of the 4 have the same slope?

OpenStudy (anonymous):

the first and last one

OpenStudy (anonymous):

No the first one has a perpendcular slope of 1/3. Which equations have -3 in front of x?

OpenStudy (anonymous):

B does

OpenStudy (anonymous):

Yeah, B and D... so now we have to find out which of those two is the correct one. Can you write the equasion where you use -1 for x and 6 for y with the same slope as the original (with -3) and then add +b.

OpenStudy (anonymous):

B would be correct bc since the - the 7 could not be right. so D would be wrong.

OpenStudy (anonymous):

Yes! Great work! Sorry, but I can't help you further at the moment, Have my own math problems to solve :-)

OpenStudy (anonymous):

lol thanxs tho a lot lol i wrote your good notes down